OpenStudy (faman39):

As incentive for Gretchen to save the 2,200 she recently inherited, her parents tell her that they will pay her 15% simple annual interest on the amount that she saves. How much money will Gretchen have after 3 years if she saves all of the money she inherited?

OpenStudy (anonymous):

Simple interest: I = p r t = 2,200 * 15% * 3 = $990 interest Thus total Gretchen saving: 2,200 + 990 = $3190